# A Neuromorphic Architecture for Scalable Event-Based Control

### v1

@@ Abstract
This paper introduces the “rebound Winner-Take-All (RWTA)” motif as the basic element of a scalable neuromorphic control architecture. From the cellular level to the system level, the resulting architecture combines the reliability of discrete computation and the tunability of continuous regulation: it inherits the discrete computation capabilities of winner-take-all state machines and the continuous tuning capabilities of excitable biophysical circuits. The proposed event-based framework addresses continuous rhythmic generation and discrete decision-making in a unified physical modeling language. We illustrate the versatility, robustness, and modularity of the architecture through the nervous system design of a snake robot.
@@ 1 Introduction
The divide between the discrete and the continuous pervades today’s control design and technology: “high-level” system design is separated from “low-level” control design, automation is separated from regulation, decision-making is separated from physical interaction. Discrete methodologies use the discrete language of finite state machines while continuous methodologies use the continuous modelling language of differential equations. The resulting need to interface discrete automata and continuous differential equations has made the theory of hybrid and cyberphysical systems a key component of control design. The advantages of separating the discrete and the continuous design are many. Yet, this separation has become a bottleneck in the design of scalable architectures that can control and regulate across a range of spatial and temporal scales [sepulchre_control_2019, matni_quantitative_2024].
@@
The divide between automation and regulation does not exist in animal machines. Nervous systems make decisions and regulate muscle activation with one and the same architecture that is event-based rather than continuous or discrete. Events occur in physical circuits obeying continuous-time differential equations, yet they can be counted and they can be modelled as the discrete events of a finite state machine [sepulchre_spiking_2022, ribar_neuromorphic_2021, schmetterling_neuromorphic_2024]. Neuromorphic control [ribar_neuromorphic_2021] aims at mixing discrete computation and continuous regulation by borrowing the physical principles and the architecture of excitable neuronal circuits. These neuromorphic approaches build upon four decades of progress in analogue and mixed-signal VLSI that have turned Carver Mead’s vision of “computing with physics” into wafer-scale neuromorphic substrates capable of emulating large-scale neural networks [mead_analog_1989, indiveri_winner-take-all_1997, indiveri_neuromorphic_2011, benjamin_neurogrid_2014, furber_spinnaker_2014, kudithipudi_neuromorphic_2025].
@@
Earlier work in neuromorphic control by the authors has focused on the control and modulation of single neurons [ribar_neuromorphic_2021] or small circuits of few neurons [schmetterling_neuromorphic_2024]. In contrast, the aim of the present paper is to introduce a scalable neuromorphic control architecture for the design of hierarchical event-based machines.
@@
The core element of our methodology is to combine the “rebound excitability” of isolated biophysical neurons with the “Winner-Take-All” paradigm of network computation. The two concepts are classical but, to the best of our knowledge, they have always been studied separately. Rebound excitability is a cellular property that has long been identified as a core mechanism of rhythm generation [brown_intrinsic_1911]. It has long been studied in neurophysiology and has inspired the design of artificial central pattern generators in rhythmic machines [ijspeert_central_2008, ijspeert_simulation_2005, matsuoka_sustained_1985, matsuoka_analysis_2011]. Winner-Take-All computation is a network property at the core of decision-making mechanisms in computational biology and in computer science [maass_computational_2000, usher_time_2001, lazzaro_winner-take-all_1988, neftci_synthesizing_2013, grossberg_adaptive_1976]. It has played a central role in the development of neuromorphic computation. A central message of this paper is to show that continuous rhythm generation and discrete decision making become one and same feature when combining the cellular property of rebound excitability and the network property of winner-take-all computation.
@@
The rest of the paper is organized as follows: Section 2 reviews the cellular property of rebound excitability in biophysical and neuromorphic circuits. Section 3 introduces the logic gates motifs and Winner-Take-All network. Section 4 introduces the rebound Winner-Take-All network and explains how different patterns can be generated with rebound Winner-Take-All network motifs. Section 5 describes how external events (pulses/biases) tune and interact with these motifs. Section 6 extends the rebound Winner-Take-All network concept from network of cells to network of smaller networks. Section 7 composes the motifs into a five-link snake controller across muscle, pattern, and supervision layers. Discussions and Conclusions are in Sections 8 and 9. All parameter details for all the plots are given in the appendix.
@@
Rhythm generation is a long standing question of neurophysiology rooted in the pioneering concept of central pattern generator (CPG) [brown_intrinsic_1911, brown_nature_1914, wilson_central_1961], whose interaction with sensory feedback [rossignol_dynamic_2006, guertin_mammalian_2009] and neuromodulation [marder_central_2001, dirk_central_2015] generates robust rhythmic activity throughout nervous systems. Computational studies explain how reciprocal inhibition and post-inhibitory rebound (PIR) produce diverse oscillatory modes in the so-called Half-Center Oscillator (HCO) [skinner_mechanisms_1994, morozova_reciprocally_2022, matsuoka_analysis_2011]. The biology of rhythms has inspired a lot of research in robotics and neuromorphic engineering, from the early adaptive CPG chip that powered an under-actuated leg [lewis_toward_2000] to recent spiking implementations on a salamander robot that can swim and walk [ijspeert_central_2008, ijspeert_simulation_2005, yu_survey_2014, lozano_control_2016], regulate a pendulum [schmetterling_neuromorphic_2024], ankle-rehabilitation devices [luo_preliminary_2024], hexapods [lopezosorio_neuromorphic_2024], compliant manipulators [abadia_neuromechanics_2025] and quadrupeds with online CPG optimization [zhang_online_2024]. Throughout this literature, the design of the controller separates system level computation tasks from CPG rhythm generation, connecting the two with ad-hoc glue logic [linares-barranco_towards_2022].
@@
Winner-Take-All (WTA) networks implement one of the core computational functions of analog computation: they perform analogue max of softmax computation and serve as the “logic gates” of continuous-time computation [majani_k-winners-take-all_1988, maass_computational_2000, oster_computation_2009, grossberg_adaptive_1976].
Early continuous-time Hopfield circuits demonstrated $k$-WTA selection [majani_k-winners-take-all_1988], and computational complexity studies confirmed the expressive power of WTA modules [maass_computational_2000]. Indiveri and colleagues translated these ideas to analogue VLSI, proposing compact soft-WTA arrays for low-power sensing and decision [indiveri_winner-take-all_1997, indiveri_artificial_2009, indiveri_neuromorphic_2011]. In spiking hardware, Oster et al. showed rapid discrimination under realistic input statistics [oster_computation_2009], while Linares-Barranco et al. integrated a WTA core with CPGs on a mixed-signal processor [linares-barranco_towards_2022]. More recently, Cotteret et al. embed finite-state machines in attractor-based spiking networks using vector-symbolic arithmetic (VSA) to configure a Hopfield weight matrix, with states encoded in rate space [cotteret_vector_2024, cotteret_distributed_2025].
Beyond hardware, a number of studies have enabled computational functions that are essential in human high-level cognitive systems to be implemented on spiking devices, such as finite state machines, auto associative memories [hopfield_neural_1982, abrahamsen_time_2004, indiveri_artificial_2009, cotteret_vector_2024, cotteret_distributed_2025] and logic gates, to be directly implemented in recurrent spiking dynamics.
@@
The concept of generating dynamical spatio-temporal patterns from winner-take-all principles also has a long history, best illustrated by the principle of winnerless competition in neural dynamics [rabinovich_dynamical_2001]. Winnerless competition circuits interconnect rate neuronal models to build heteroclinic trajectories between saddle points, which raises challenges in ensuring both sensitivity to inputs and robustness against noise [rabinovich_dynamical_2001, afraimovich_heteroclinic_2004]. A number of studies try to resolve this issue by determining the condition of robust heteroclinic [rabinovich_discrete_2018, bick_dynamical_2009, seliger_dynamics-based_2003, ashwin_designing_2013, rabinovich_robust_2014, ashwin_criteria_2011] or adding external controllers [herron_robust_2023]. Previous work has shown that winnerless competition controllers deliver tunable patterned motion, including snake locomotion [horchler_designing_2015].
@@ 2.1 Rebound Excitability
The mechanism of rebound excitability has been studied in great detail since the early days of neurophysiology. We refer the reader to [ribar_neuromorphic_2021] for a detailed survey of the relevant physiology as well as the neuromorphic question of designing artificial neurons with tunable rebound excitability, we refer the reader to [ribar_neuromorphic_2021].
In this paper, we rely on the classical circuit architecture laid out in the pioneering work of Hodgkin and Huxley [hodgkin_quantitative_1952]. A neuron is modeled as an electrical circuit that connects one leaky capacitor (the passive membrane) to a parallel bank of active filters defined by the series combination of a battery and a memristive element, as shown in Figure 2. Each memristive element obeys the Ohmic law $i=gv$, where the ‘conductance’ $g$ depends nonlinearly and dynamically on the membrane voltage.
@@
Excitability is the key property of a biophysical neuron [sepulchre_excitable_2018]: the voltage response to a current stimulus is passive for small stimuli and event-based when the stimulus exceeds a threshold. The simplest event-based response of a neuron is a single spike. An important property of biophysical neurons is that both depolarizing inputs (excitation - the input current flow increases the membrane voltage) and hyperpolarizing inputs (inhibition - the input current flow decreases the membrane voltage) can trigger an event. In the latter case, the event occurs at the end of the inhibition stimulus. Such an event, illustrated in Figure 3(a), is called a rebound event, and the corresponding excitability property is called rebound excitability.
@@ 2.2 Rebound Spiking and Bursting
Throughout the paper, we focus on qualitative properties that do not depend on fine details of conductance models. Therefore, for simplicity, we adopt the model proposed in [ribar_neuromorphic_2021]. This model captures the essential properties of neurons, including rebound excitability.
@@ 2.3 Inhibitory and Excitatory Synapses
Synaptic currents model directional interconnections between neurons. A synaptic current is (dynamically) determined by the presynaptic neuron membrane voltage and affects the postsynaptic neuron membrane voltage. In this paper we adopt the simplified synaptic model of [ribar_neuromorphic_2021]. A synaptic current $I_{ij}$ of type $k$ from the presynaptic neuron $j$ (voltage $V_{j}$) to the postsynaptic neuron $i$ has the expression
@@ 3 Functional Level: Decision Making
Logical operations such as NOT, AND and OR are the fundamental elements of decision making. In this section, we provide a neuromorphic realization of these classical logical elements. We then move to a more complex network, illustrating a neuromorphic realization of the classical Winner-Take-All mechanisms (WTA) [grossberg_adaptive_1976]. Logic gates and winner-take-all mechanism realizations are illustrated in Figure 5.
@@ 4 Functional Level: Pattern Generation
It consists of two rebound neurons reciprocally connected by inhibitory synapses, as shown at the bottom row of Figure 7. It is the simplest pattern generator built from rebound neurons [brown_intrinsic_1911].
A useful way to understand the network is to start with a single inhibitory synapse from neuron 1 to neuron 2 (Figure 7, top row). When neuron 1 spikes, it inhibits neuron 2; upon release of inhibition, neuron 2 produces a rebound spike. With reciprocal inhibition, this rebound triggers the reverse action, yielding an alternating left–right pattern. This rebound–based mechanism is classical and extensively documented [marder_central_2001, dirk_central_2015].
@@
Spiking Ring Oscillator: A ring of $N$ spiking neurons with fast all–to–all inhibition and slow excitation connections forming a ring produces an autonomous traveling wave $1\!\to\!2\!\to\!\cdots\!\to\!N$, in the absence of external input. The length of the inter spike interval is given by the combined contribution of spike interval and rebound interval (see Figure 3(a)). The former is mostly affected by the time constant of the fast inhibitory synapses. The latter is sensitive to the time constant and weight of the slow excitatory synapses and it also increases monotonically with the input magnitude. More details can be found in [huo_winner-takes-all_2025].
@@
We complete this section by emphasizing that the computational principles of our neuromorphic architecture are consistent with the biophysical architecture of EI networks studied in neuroscience [wilson_excitatory_1972, wilson_spikes_1999]. An EI network of rebound neurons was previously studied in [drion_switchable_2018] to show how cellular neuromodulation can easily reconfigure the network states. A neuromorphic neuron in our architecture corresponds to a pair of excitatory and inhibitory neurons in EI networks. The inhibitory layer must be densely connected to approximate the all-to-all topology of a rebound network, and the inhibitory synapses are typically of the GABA type. In contrast, the excitatory layer is sparsely connected by AMPA synapses. While AMPA synapses are normally faster than GABA synapses, our slow excitatory synapse can be regarded as a model of the synaptic delay resulting from event propagation through several nodes in the excitatory layer.
@@ 7.1 Snake mechanics, muscle actuation, and neuromorphic controller basics
The plant: the snake ’skeleton’ consists of five rigid segments connected by four revolute joints. Each joint is actuated by a pair of ‘muscles’. Each muscle is modeled by (i) a torque spring with a variable reference point, (ii) a linear torque spring, and (iii) a linear damper [mirbagheri_intrinsic_2000, feldman_once_1986], as shown in Figure 11. The contraction/relax of a muscle is modeled via the change in the reference trajectory of the first spring. In what follows, we regard the combination of snake mechanics and muscles as the plant.
@@ 8 Discussion
Figure 21 can be regarded as a revision of the traditional design layout in Figure 1 in the event-based language of the present paper. A key feature of the rebound Winner-Take-All architecture is that all the cells in Figure 21 use the same modeling language and the same set of interconnection rules. Design, tuning, and adaptation are addressed at any level and any layer with the same concept of synaptically interconnected rebound neurons.
One methodology across layers and levels At the event level, all signals (sensory, internal, motor) are represented as spike or burst events generated by rebound-excitable neurons. At the functional level, the same rebound Winner-Take-All architecture implements both the rhythmic functionalities of pattern generation and the logical gates of an automaton. At the system level, sub-networks become states of a higher level automaton obeying the same architecture.
The unified language is in contrast to the traditional separation between the discrete language of automata for “higher level” design and the continuous (physical) language of differential equations for “lower level” design. Decision-making and regulation are merged in the same event-based framework, allowing continuous tuning and discrete topological/functional design at each scale.
@@
Unified tuning. Because all layers and levels share the same substrate, they share the same tuning properties: (i) Changing neuron internal parameters to tune the shape and duration of an event (spike or burst) (ii) common or neuron-specific bias currents, which continuously tune the rebound time and thus rhythm frequency (demonstrated in Figure 9); (iii) transient sensory pulses, which apply immediate phase selection or hard resets by temporally biasing the current WTA competition. These three mechanisms are sufficient to modulate behaviours at every layer without introducing layer-specific control laws. As a result, adaptation is uniform: the same type of parameter change (bias or synaptic time constant/gain) has predictable, monotone effects from lower layer (muscle actuation) to higher layer (supervisory control).
@@
How this differs from existing architectures. Conventional neuromorphic locomotion task typically couple (i) a continuous CPG for rhythms, (ii) a discrete finite-state machine for mode logic, and (iii) separate low-level joint controllers, glued together with cross-domain interfaces [lozano_control_2016, linares-barranco_towards_2022, zhang_online_2024]. This heterogeneity forces different design languages (ODEs for regulation, automata or associative memories for logic [cotteret_vector_2024, cotteret_distributed_2025]), incurs interface fragility, and complicates tuning (parameters have layer-specific meanings). In contrast, the proposes framework avoids the need for ad-hoc glue logic between rate-based and event-based blocks, and no change of modeling paradigm between decision making and trajectory planning.

### v2

@@ Abstract
This paper introduces the “rebound Winner-Take-All (RWTA)” motif as the basic element of a scalable neuromorphic control architecture. From the cellular level to the system level, the resulting architecture combines the reliability of discrete computation and the tunability of continuous regulation: it inherits the discrete computation capabilities of winner-take-all state machines and the continuous tuning capabilities of excitable biophysical circuits. The proposed event-based framework addresses continuous rhythmic generation and discrete decision-making in a unified physical modelling language. We illustrate the versatility, robustness, and modularity of the architecture through the nervous system design of a snake robot.
@@ 1 Introduction
The divide between the discrete and the continuous pervades today’s control design and technology: “high-level” system design is separated from “low-level” control design, automation is separated from regulation, decision-making is separated from physical interaction. Discrete methodologies use the discrete language of finite state machines while continuous methodologies use the continuous modelling language of differential equations. The resulting need to interface discrete automata and continuous differential equations has made the theory of hybrid and cyberphysical systems a key component of control design. The advantages of separating the discrete and the continuous design are many. Yet, this separation has become a bottleneck in the design of scalable architectures that can control and regulate across a range of spatial and temporal scales [61, 44].
@@
The divide between automation and regulation does not exist in animal machines. Nervous systems make decisions and regulate muscle activation with one and the same architecture that is event-based rather than continuous or discrete. Events occur in physical circuits obeying continuous-time differential equations, yet they can be counted and they can be modelled as the discrete events of a finite state machine [63, 55, 59]. Neuromorphic control [55] aims at mixing discrete computation and continuous regulation by borrowing the physical principles and the architecture of excitable neuronal circuits. These neuromorphic approaches build upon four decades of progress in analogue and mixed-signal VLSI that have turned Carver Mead’s vision of “computing with physics” into wafer-scale neuromorphic substrates capable of emulating large-scale neural networks [47, 31, 30, 7, 18, 33].
@@
Earlier work in neuromorphic control by the authors has focused on the control and modulation of single neurons [55] or small circuits of few neurons [59]. In contrast, the aim of the present paper is to introduce a scalable neuromorphic control architecture for the design of hierarchical event-based machines.
@@
The core element of our methodology is to combine the “rebound excitability” of isolated biophysical neurons with the “Winner-Take-All” paradigm of network computation. The two concepts are classical but, to the best of our knowledge, they have always been studied separately. Rebound excitability is a cellular property that has long been identified as a core mechanism of rhythm generation [10]. It has long been studied in neurophysiology and has inspired the design of artificial central pattern generators in rhythmic machines [28, 27, 45, 46]. Winner-Take-All computation is a network property at the core of decision-making mechanisms in computational biology and in computer science [40, 65, 34, 50, 19]. It has played a central role in the development of neuromorphic computation. A central message of this paper is to show that continuous rhythm generation and discrete decision making become one and same feature when combining the cellular property of rebound excitability and the network property of winner-take-all computation.
@@
The rest of the paper is organized as follows: Section 2 reviews the cellular property of rebound excitability in biophysical and neuromorphic circuits. Section 3 introduces the logic gates motifs and Winner-Take-All network. Section 4 introduces the rebound Winner-Take-All network and explains how different patterns can be generated with rebound Winner-Take-All network motifs. Section 5 describes how external events (pulses/biases) tune and interact with these motifs. Section 6 extends the rebound Winner-Take-All network concept from network of cells to network of smaller networks. Section 7 composes the motifs into a five-link snake controller across muscle, pattern, and supervision layers. Discussions and Conclusions are in Sections 8 and 9. All parameter details for all the plots are given in the appendix. Related code can be found in https://github.com/Huoyongtony/A-Neuromorphic-Architecture-for-Scalable-Event-Based-Control.
@@
Rhythm generation is a long standing question of neurophysiology rooted in the pioneering concept of central pattern generator (CPG) [10, 9, 67], whose interaction with sensory feedback [56, 20] and neuromodulation [42, 14] generates robust rhythmic activity throughout nervous systems. Computational studies explain how reciprocal inhibition and post-inhibitory rebound (PIR) produce diverse oscillatory modes in the so-called Half-Center Oscillator (HCO) [64, 49, 46]. The biology of rhythms has inspired a lot of research in robotics and neuromorphic engineering, from the early adaptive CPG chip that powered an under-actuated leg [35] to recent spiking implementations on a salamander robot that can swim and walk [28, 27, 70, 38], regulate a pendulum [59], ankle-rehabilitation devices [39], hexapods [37], compliant manipulators [1] and quadrupeds with online CPG optimization [71]. Throughout this literature, the design of the controller separates system level computation tasks from CPG rhythm generation, connecting the two with ad-hoc glue logic [36].
@@
Winner-Take-All (WTA) networks implement one of the core computational functions of analog computation: they perform analogue max of softmax computation and serve as the “logic gates” of continuous-time computation [41, 40, 51, 19].
Early continuous-time Hopfield circuits demonstrated $k$-WTA selection [41], and computational complexity studies confirmed the expressive power of WTA modules [40]. Indiveri and colleagues translated these ideas to analogue VLSI, proposing compact soft-WTA arrays for low-power sensing and decision [31, 29, 30]. In spiking hardware, Oster et al. showed rapid discrimination under realistic input statistics [51], while Linares-Barranco et al. integrated a WTA core with CPGs on a mixed-signal processor [36]. More recently, Cotteret et al. embed finite-state machines in attractor-based spiking networks using vector-symbolic arithmetic (VSA) to configure a Hopfield weight matrix, with states encoded in rate space [13, 12].
Beyond hardware, a number of studies have enabled computational functions that are essential in human high-level cognitive systems to be implemented on spiking devices, such as finite state machines, auto associative memories [24, 2, 29, 13, 12] and logic gates, to be directly implemented in recurrent spiking dynamics.
@@
The concept of generating dynamical spatio-temporal patterns from winner-take-all principles also has a long history, best illustrated by the principle of winnerless competition in neural dynamics [52]. Winnerless competition circuits interconnect rate neuronal models to build heteroclinic trajectories between saddle points, which raises challenges in ensuring both sensitivity to inputs and robustness against noise [52, 3]. A number of studies try to resolve this issue by determining the condition of robust heteroclinic [54, 8, 60, 5, 53, 4] or adding external controllers [22]. Previous work has shown that winnerless competition controllers deliver tunable patterned motion, including snake locomotion [25].
@@ 2.1 Rebound Excitability
The mechanism of rebound excitability has been studied in great detail since the early days of neurophysiology. We refer the reader to [55] for a detailed survey of the relevant physiology as well as the neuromorphic question of designing artificial neurons with tunable rebound excitability, we refer the reader to [55].
In this paper, we rely on the classical circuit architecture laid out in the pioneering work of Hodgkin and Huxley [23]. A neuron is modeled as an electrical circuit that connects one leaky capacitor (the passive membrane) to a parallel bank of active filters defined by the series combination of a battery and a memristive element, as shown in Figure 2. Each memristive element obeys the Ohmic law $i=gv$, where the ‘conductance’ $g$ depends nonlinearly and dynamically on the membrane voltage.
@@
Excitability is the key property of a biophysical neuron [62]: the voltage response to a current stimulus is passive for small stimuli and event-based when the stimulus exceeds a threshold. The simplest event-based response of a neuron is a single spike. An important property of biophysical neurons is that both depolarizing inputs (excitation - the input current flow increases the membrane voltage) and hyperpolarizing inputs (inhibition - the input current flow decreases the membrane voltage) can trigger an event. In the latter case, the event occurs at the end of the inhibition stimulus. Such an event, illustrated in Figure 3(a), is called a rebound event, and the corresponding excitability property is called rebound excitability.
@@ 2.2 Rebound Spiking and Bursting
Throughout the paper, we focus on qualitative properties that do not depend on fine details of conductance models. Therefore, for simplicity, we adopt the model proposed in [55]. This model captures the essential properties of neurons, including rebound excitability.
@@ 2.3 Inhibitory and Excitatory Synapses
Synaptic currents model directional interconnections between neurons. A synaptic current is (dynamically) determined by the presynaptic neuron membrane voltage and affects the postsynaptic neuron membrane voltage. In this paper we adopt the simplified synaptic model of [55]. A synaptic current $I_{ij}$ of type $k$ from the presynaptic neuron $j$ (voltage $V_{j}$) to the postsynaptic neuron $i$ has the expression
@@ 3 Functional Level: Decision Making
Logical operations such as NOT, AND and OR are the fundamental elements of decision making. In this section, we provide a neuromorphic realization of these classical logical elements. We then move to a more complex network, illustrating a neuromorphic realization of the classical Winner-Take-All mechanisms (WTA) [19]. Logic gates and winner-take-all mechanism realizations are illustrated in Figure 5.
@@ 4 Functional Level: Pattern Generation
It consists of two rebound neurons reciprocally connected by inhibitory synapses, as shown at the bottom row of Figure 7. It is the simplest pattern generator built from rebound neurons [10].
A useful way to understand the network is to start with a single inhibitory synapse from neuron 1 to neuron 2 (Figure 7, top row). When neuron 1 spikes, it inhibits neuron 2; upon release of inhibition, neuron 2 produces a rebound spike. With reciprocal inhibition, this rebound triggers the reverse action, yielding an alternating left–right pattern. This rebound–based mechanism is classical and extensively documented [42, 14].
@@
Spiking Ring Oscillator: A ring of $N$ spiking neurons with fast all–to–all inhibition and slow excitation connections forming a ring produces an autonomous traveling wave $1\!\to\!2\!\to\!\cdots\!\to\!N$, in the absence of external input. The length of the inter spike interval is given by the combined contribution of spike interval and rebound interval (see Figure 3(a)). The former is mostly affected by the time constant of the fast inhibitory synapses. The latter is sensitive to the time constant and weight of the slow excitatory synapses and it also increases monotonically with the input magnitude. More details can be found in [26].
@@
We complete this section by emphasizing that the computational principles of our neuromorphic architecture are consistent with the biophysical architecture of EI networks studied in neuroscience [69, 68]. An EI network of rebound neurons was previously studied in [15] to show how cellular neuromodulation can easily reconfigure the network states. A neuromorphic neuron in our architecture corresponds to a pair of excitatory and inhibitory neurons in EI networks. The inhibitory layer must be densely connected to approximate the all-to-all topology of a rebound network, and the inhibitory synapses are typically of the GABA type. In contrast, the excitatory layer is sparsely connected by AMPA synapses. While AMPA synapses are normally faster than GABA synapses, our slow excitatory synapse can be regarded as a model of the synaptic delay resulting from event propagation through several nodes in the excitatory layer.
@@ 7.1 Snake mechanics, muscle actuation, and neuromorphic controller basics
The plant: the snake ’skeleton’ consists of five rigid segments connected by four revolute joints. Each joint is actuated by a pair of ‘muscles’. Each muscle is modeled by (i) a torque spring with a variable reference point, (ii) a linear torque spring, and (iii) a linear damper [48, 16], as shown in Figure 11. The contraction/relax of a muscle is modeled via the change in the reference trajectory of the first spring. In what follows, we regard the combination of snake mechanics and muscles as the plant.
@@ 8.1 Hierarchical event design
Figure 21 can be regarded as a revision of the traditional design layout in Figure 1 in the event-based language of the present paper. A key feature of the rebound Winner-Take-All architecture is that all the cells in Figure 21 use the same modelling language and the same set of interconnection rules. Design, tuning, and adaptation are addressed at any level and any layer with the same concept of synaptically interconnected rebound neurons.
The central ingredients that make this architectural notion of scale concrete are (i) the separation of time-scales across the hierarchy and (ii) the separation of cellular excitability (event generation) from WTA-style competition (event orchestration). At the event level, all signals (sensory, internal, motor) are represented as spike or burst events generated by rebound-excitable neurons. Importantly, the semantics of an “event” evolve with scale: at the lowest level, events are spikes/bursts that encode changes in a physical variable or trigger a local actuation/entrainment mechanism; at higher levels, events can be interpreted as symbolic or supervisory decisions (e.g., “enable forward wave”, “enable right-turn strategy”).
This unification is expressed not only across different representations but also across different functions. The same rebound Winner-Take-All architecture can implement the rhythmic functions of pattern generation and the logical functions of an automaton. Higher level events consist of specific sequences of lower level events, obeying the same modelling principles. This hierarchical event composition is what enables an end-to-end event-based representation that resembles the layered structure of conventional hiearchical control while remaining confined within one modeling language across scales.
@@
Because all layers and levels share the same modeling language, they also share the same tuning properties: (i) changing neuron internal parameters to tune the shape and duration of an event (spike or burst); (ii) bias currents continuously tune the rebound time and thus rhythm frequency (demonstrated in Figure 9); (iii) sensory pulses, which apply immediate phase selection or hard resets by temporally biasing the current WTA competition. These three mechanisms are sufficient to modulate behaviours at every layer without introducing layer-specific control laws. As a result, adaptation is uniform: the same type of parameter change (bias or synaptic time constant/gain) has predictable, monotone effects from lower layer (muscle actuation) to higher layer (supervisory control).
@@
Finally, it is worth emphasizing how this differs from existing architectures. Conventional neuromorphic locomotion tasks typically couple (i) a continuous CPG for rhythms, (ii) a discrete finite-state machine for mode logic, and (iii) separate low-level joint controllers, glued together with cross-domain interfaces [38, 36, 71]. This heterogeneity forces different design languages (ODEs for regulation, automata or associative memories for logic [13, 12]), incurs interface fragility, and complicates tuning (parameters have layer-specific meanings). In contrast, the proposed framework avoids the need for ad-hoc glue logic between rate-based and event-based blocks, and avoids a change of modelling paradigm between decision making and trajectory planning. The end-to-end event-based representation makes the supervisory layer naturally interpretable as event-driven mode selection, which also invites connections to hybrid and variable-structure viewpoints (e.g., switching surfaces and mode logic), while still grounding the implementation in the same excitable physical primitives.
8.2 A neuromorphic architecture…
The central feature of the architecture proposed in this paper is that it is rooted in the physiology of neuronal excitability, and, singularly, rebound excitability. To the best of the authors knowledge, it is the first proposed event-based design architecture of this type, and it makes it eminently neuromorphic. Neuronal excitability is a core feature of any nervous system, but it has not been embraced by artificial intelligence technology. Our research suggests the potential of neuronal excitability for neuromorphic engineering:
  1. 1.
    The design of an event-based behavior requires the design of events and the design of sequences of events. Neuronal excitability is the property that enables a complete decoupling between the two: shaping the event is achieved by regulating internal conductances, whereas shaping the sequence of events is achieved by regulating external conductances. The hierarchy of our architecture comes from the fact that events at coarser scales are made of sequences of events at finer scales. For this reason, the decoupling between internal and external repeats itself across scales, with an evolving semantics for the functionality of events: spike control regulates burst events, burst control regulates pattern events, pattern control regulates locomotion events, locomotion control regulates behavior, etc..
  2. 2.
    The unique property of excitability to regulate across scales has been advocated in neurophysiology [43], in computational neuroscience [6], and in control [61]. Our architecture is consistent with those principles. At every scale, an event is modeled by a neuromorphic circuit with the same conductance architecture, enabling regulation at every scale from the same adaptive mechanisms [11, 58]. At every scale, the topology of the network constrains the sequence of events from the same complementary roles of excitatory and inhibitory synaptic interconnections. Those principles seem aligned with those long advocated in biological cybernetics [68].
  3. 3.
    Excitability seems also unique in making heterogeneity a design feature rather than a design nuisance: events are tuned independently from each other, de facto resulting in populations of heterogenous events at any scale. In turn, this very heterogeneity becomes a tuning feature for the sequence of events that will define the event at the next scale. This feature was illustrated at an elementary level by tuning the period of a ring oscillator by inserting a burster in a ring of spiking neurons. Again, this concrete hierarchical tuning procedure seems consistent with recent studies on the value of heterogeneity in tuning and learning [21].
8.3 … for scalable event-based control
There is a long way from designing a scalable architecture to designing a scalable machine. The present paper should only be regarded as a first and conceptual step towards scalable event-based control. We envision however that the scalability of the control design will benefit from the scalability of the architecture. In particular, the following three distinct avenues provide encouraging research directions:
  1. 1.
    Modeling, analysis, and synthesis at scale: a hierarchical architecture calls for hierarchical modeling, analysis, and design tools. Shaping an event at a coarser scale must be abstracted from shaping the event at a finer scale. The article [57] is a preliminary step in that direction: a hierarchy of discrete event models is extracted from the same continuous neuromorphic system to enable tractability across scales from tractability at scale.
  2. 2.
    Tuning and learning at scale: classical principles from adaptive control are sufficient to tune at cellular scale [11, 58]. Currently, those principles do not extend across scales. We envision however that the same principles could be implemented at any scale, pending a systematic methodology to model the neuromorphic system at different scales. Similarly, classical Hopfield learning paradigms seem applicable at one scale (for instance, to learn a particular sequence from a all-to-all inhibitory architecture), and we envision to apply those same learning principles at any scale. This particular research direction will be the topic of a forthcoming publication by the authors.
  3. 3.
    Physical realisations at scale. The question of turning the conceptual architecture into a physical realisation is not addressed in this paper but it deserves particular attention. The proposed architecture is an interconnection of memristive elements. At each scale, the design of new events requires memristive elements with commensurate temporal and spatial scales: spikes require fast conductances, bursts require slower conductances, sequences of bursts require even slower synapses, and so on.
    The same hierarchy is found in the current research on memristive materials. Each physical domain comes with its own physical scales. It is a challenge to cover many scales in one specific domain, but much promise arises from interconnecting memristive devices across different physical domains (electronics, organics, pneumatics,…), and meristive designs are investigated in electronics, pneumatics, organics. Encouraging steps have been taken in implementing rebound bursting neurons on neuromorphic chips [32, 66]. Research labs are currently exploring similar realisations in the pneumatic and organic domain. Our hierarchical architecture is well suited to exploring multi-scale implementations by exploiting multi-scale physical domains.
    Besides physical realisations of neuromorphic architectures, more and more attention is also devoted to digital realisations of neuromorphic architectures. A digital implementation of the proposed neuromorphic architecture on dedicated chips such as [17] is a most promising avenue in the near future.